OpenStudy (luckycoins888):

fan + medal!! Convert the decimal to a percent. 0.007 A. 0.007% B. 0.07% C. 0.7% D. 7%

OpenStudy (luckycoins888):

@quentavess @Walkerman15 @whysignupfor1question @eliassaab @RainbowsxD @RedNeckOutLaw @redredred123 @ro561man @russano @Romantico @Taylor<3sRin @this_emo_chic1023 @Taylev105 @TeenWolfGirl @TheMysteriousRainbowPegasus @Thatonegirl_

OpenStudy (taylev105):

ok i can do it

OpenStudy (anonymous):

move the decimal over twice right

OpenStudy (anonymous):

0.007 * 100 = 0.7

OpenStudy (taylev105):

you are correct @quentavess

sammixboo (sammixboo):

All you need to do is multiply the decimal by 100 and then add the percent sign 0.08 = 8% 0.08*100 = 8 0.000009 = 0.0009% 0.000009*100 = 0.0009 0.4 = 40% 0.4*100 = 40

OpenStudy (thatonegirl_):

To go from a percent to decimal, divide by 100 or move the decimal point to the left 2. To go from a decimal to percent, multiply by 100 or move the decimal point to the right 2.